OpenStudy (anonymous):

Given the function f(x) = 6(x+2) − 3, solve for the inverse function when x = 21.

OpenStudy (anonymous):

for the inverse you're going to switch x and y so your new equation will be x=6(y+2)-3 and then we will solve for x by first adding 3 over so x+3=6(y+2) and then divide by 6 so ((x+3)/6)=y+2 and finally subtract that 2 over leaving you with \[\frac{ (x+3) }{ 6 } -2 = y\] then plug in 21 for x so 21+3 = 24, divide that by 6 and you get 4 and then subtract 2 and you get 2! lemme know if you need more help

OpenStudy (anonymous):

Thank you! Could you help me with another question?

OpenStudy (anonymous):

yes ma'am throw it up here

OpenStudy (anonymous):

Which function below is the inverse of f(x) = x2 − 25?

OpenStudy (anonymous):

so I'm gonna make you work a little more on this one but I'll support you through it. okay so the inverse means we will switch x and y so we will have x=y^2 - 25 and were trying to solve for y right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so we would go about doing that by first moving the 25 over to the other side by adding it and then to solve for y we must take the square root of the entire other side

OpenStudy (anonymous):

so it would be the square root of x+25=y